OpenStudy (anonymous):

Which of the following is the equation of a line that passes through the points (0,6) and (2,10)?

OpenStudy (anonymous):

To find slope: (10-6)/(2-0) m=2 y-6=2(x-0) y-6=2x y=2x+6